OpenStudy (anonymous):

HELP URGENT! EXAM AFTER 3 HOURS, INTEGRATION

OpenStudy (anonymous):

can anybody explain me how to find volume integration in short?

OpenStudy (anonymous):

Question Attached

OpenStudy (dumbcow):

 \[V = \pi \int\limits_{-2}^{2}R^{2}(x) - r^{2} (x) dx\] \[R(x) = (9-x^{2}) +1\] \[r(x) = (x^{2} +1) +1\]